# Distributed Primal-Dual Algorithms: Unification, Connections, and Insights

### v1

@@ 2.3 Distributed Proximal ADMM
Deng & Yin (2016) introduces an additional proximal term into the standard ADMM algorithm for solving $\min_{x,y}f(x)+g(y)$ subject to $Ax+By=b$ , which is referred to as generalized ADMM or proximal ADMM. Applying Algorithm 2 of Deng & Yin (2016), we can solve the dual problem (2) with following update rule:
$\begin{split}v^{(t+1)}&=\underset{v\in\mathbb{R}^{n}}{\arg\min}\;\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{P}_{k}}\ell^{*}_{i}(v_{i})-\langle w^{(t)},\frac{1}{n}Xv+u^{(t)}\rangle\\ &+\frac{\rho}{2}\Big{\|}\frac{1}{n}Xv+u^{(t)}\Big{\|}^{2}+\frac{1}{2}\|v-v^{(t)}\|_{Q}^{2},\\ u^{(t+1)}&=\underset{u\in\mathbb{R}^{d}}{\arg\min}\;g^{*}(u)-\langle w^{(t)},\frac{1}{n}Xv^{(t+1)}+u\rangle\\ &+\frac{\rho}{2}\Big{\|}\frac{1}{n}Xv^{(t+1)}+u\Big{\|}^{2},\\ w^{(t+1)}&=w^{(t)}-\rho\left(\frac{1}{n}Xv^{(t+1)}+u^{(t+1)}\right),\end{split}$
where $\rho>0$ is the tuning parameter and $Q$ is a positive semi-definite matrix.
Motivated by the Deng et al. (2017) that solves distributed optimization problems without the regularization term, we identified two choices of the positive semi-definite matrices $Q$s based on the data matrix $X_{[k]}$s to solve (2), which enables parallel updates of $v$ across $K$ machines:
$Q_{1}=\frac{\rho}{n^{2}}(\eta_{1}\,\text{diag}(X_{[1]}^{\top}X_{[1]},\ldots,X_{[K]}^{\top}X_{[K]})-X^{\top}X);$
$Q_{2}=\frac{\rho}{n^{2}}\left(\eta_{2}I-X^{\top}X\right)$
Here $\eta_{1}$ and $\eta_{2}$ are chosen to ensure that $Q_{1}$ and $Q_{2}$ are positive semi-definite. The resulted distributed proximal ADMM algorithms can both be cast into a compact update form that involves only $w$-steps and $v$-steps.
Proposition 2.2.
The updating formula (6) and (7) below respectively result in the same iterative sequence of $\{v^{(t)}\}$ and $\{w^{(t)}\}$ as the distributed proximal ADMM algorithm with choices of matrices $Q_{1}$ and $Q_{2}$.
$v^{(t+1)}_{[k]}=$ $\underset{v_{[k]}\in\mathbb{R}^{n_{k}}}{\arg\min}$
$\frac{1}{n}\sum_{i\in\mathcal{P}_{k}}\ell^{*}_{i}(v_{i})+\frac{\rho\eta_{1}}{2n^{2}}\Big{\|}v_{[k]}-v_{[k]}^{(t)}\Big{\|}_{X_{[k]}^{\top}X_{[k]}}^{2}$
$-\frac{1}{n}\Big{\langle}X_{[k]}^{\top}\left(2w^{(t)}-w^{(t-1)}\right),v_{[k]}\Big{\rangle},k\in[K],$
$w^{(t+1)}=$ $\mathrm{prox}_{\rho g}\left(w^{(t)}-\frac{\rho}{n}Xv^{(t+1)}\right);$ (6)
$v_{[k]}^{(t+1)}=$
$\mathrm{prox}_{(n/\rho\eta_{2})\ell^{*}_{[k]}}\left(v_{[k]}^{(t)}+\frac{n}{\rho\eta_{2}}X_{[k]}^{\top}\left(2w^{(t)}-w^{(t-1)}\right)\right),\\ k\in[K],$
$w^{(t+1)}=\mathrm{prox}_{\rho g}\left(w^{(t)}-\frac{\rho}{n}Xv^{(t+1)}\right).$ (7)
The update rule of (6) and (7) are named Proximal-1-PD and Proximal-2-PD, respectively. The distributed proximal ADMM algorithms of either $Q$ are guaranteed to converge for any $\rho>0$. The following lemma provides a reliable choice for selecting the tuning parameters $\eta_{1}$ and $\eta_{2}$ that ensures $Q_{1}$ and $Q_{2}$ to be positive semidefinite, satisfying the requirement of Deng & Yin (2016).
Lemma 2.3.
@@
The minimal $\eta_{2}$ to let $Q_{2}\succeq 0$ is $\lambda_{\max}\left(X^{\top}X\right)$. However, this choice is practically infeasible in a federated learning setup, as it requires the aggregation of samples from all machines.
@@ 2.4 Summary
The five algorithms CoCoA-PD, Consensus-PD, LinConsensus-PD, Proximal-1-PD, and Proximal-2-PD (3)–(7) have a unified form of iterative primal and dual updates, which enables us to investigate the connection between algorithms and apply a unified analysis of convergence for these algorithms in Section 4. The updates of $v_{[k]}$ in all algorithms involve applying $\mathrm{prox}_{\ell_{[k]}^{*}}(\cdot)$ on a linear combination of $w^{(t)}$ and $v_{[k]}^{(t)}$. All updates of $w$ involve a linear combination of current $w^{(t)}$, the information transmitted from individual machines for the current time step $X_{[k]}v_{[k]}^{(t+1)}$, and the information from the previous time step $X_{[k]}v_{[k]}^{(t)}$.
An immediate benefit of using these unified primal-dual update formula, instead of their original algorithms, is that they enable a quick evaluation of the dual gaps that bound the error in objectives. Such dual gaps are computed from plugging in the updated values of $\{v_{[k]}^{(t)}\}_{k=1}^{K},w^{(t)}$ into the primal objective (1) and dual objective (2) respectively.
@@
We summarize the selection of tuning parameters in the five algorithms mentioned above. With fixed optimal parameter $\sigma=K$ and $\gamma=1$ in CoCoA algorithm (Smith et al., 2018), CoCoA-PD does not have tuning parameters. The optimal selection of parameters $\eta_{1},\eta_{2}$ in Proximal-1-PD, Proximal-2-PD, and $\tau$ in LinConsensus-PD are given in this article and confirmed by the experiments. The step sizes $\beta$ of Consensus-PD and LinConsensus-PD, and the step size $\rho$ of the Proximal-1-PD and Proximal-2-PD significantly affect the convergence speed (Boyd et al., 2011) and should be tuned in a case-specific manner, as validated in our experiments (See Section 5.1).
3 Connections Between the Primal-Dual Algorithms
@@
Corollary 3.1.
Assume that $g(w)=\lambda\left\lVert w\right\rVert_{2}^{2}$, $\rho=\lambda^{-1}$, $\gamma=1$ and $\eta_{1}=\sigma$. The update procedure of (3) and (6) will give identical updates for the dual variable $v_{[k]}^{(t)},k\in[K]$, given that both algorithm starts at the same $v_{[k]}^{(1)},k\in[K]$.
@@
It is also interesting to see if the connection between CoCoA-PD and Prixmal-1-PD can be extended to other CoCoA variants with general penalty $g$ (Smith et al., 2018). To this question, we give a negative answer, because the connection in Corollary 3.1 relies on the same quadratic structure of the ridge penalty in CoCoA and the augmented Lagrangian in the Proximal ADMM algorithm.
@@
Proximal ADMM and Consensus ADMM.
To understand the relationship between distributed Proximal ADMM and Consensus ADMM with regularization, we express the primal problem (1) into the saddle point problem:
$\begin{split}\underset{w\in\mathbb{R}^{d}}{\min}\;\underset{v\in\mathbb{R}^{n}}{\max}\;L(w;v),\end{split}$ (8)
where $L(w;v):=\frac{1}{n}\sum_{i=1}^{n}\left(v_{i}\langle w,x_{i}\rangle-\ell_{i}^{*}(v_{i})\right)+g(w)$, given that $\ell_{i}(w^{\top}x_{i})=\underset{v_{i}\in\mathbb{R}}{\sup}\left\{v_{i}\langle w,x_{i}\rangle-\ell_{i}^{*}(v_{i})\right\}$. The following lemma shows that all algorithms (4)–(7) can be represented by a generic update rules with
$P(z^{(t)}-z^{(t+1)})\in\mathcal{F}(z^{(t+1)}),$ (9)
where $z=(w,v)$ denotes the primal and dual updates, $\mathcal{F}(z)=\mathcal{F}(w,v)=\begin{pmatrix}\partial_{w}L(w;v)\\ -\partial_{v}L(w;v)\end{pmatrix}$, and $P\in\mathbb{R}^{(n+p)\times(n+p)}$ is a positive semi-definite matrix.
Proposition 3.2.
The generic update rules (9) with $P$ selected as the matrices $P_{1}$, $P_{2}$, $P_{3}$ and $P_{4}$ are the same as the update rule (4), (5), (6) and (7), corresponding to Consensus-PD, LinConsensus-PD, Proximal-1-PD, and Proximal-2-PD, respectively:
$P_{1}=\begin{pmatrix}\beta KI&\frac{1}{n}A\\ \frac{1}{n}A^{\top}&\frac{\beta^{-1}}{n^{2}}B\end{pmatrix},P_{2}=\begin{pmatrix}\beta KI&\frac{1}{n}A\\ \frac{1}{n}A^{\top}&\frac{\tau\beta^{-1}}{n^{2}}I\end{pmatrix}$
$P_{3}=\begin{pmatrix}\rho^{-1}I&-\frac{1}{n}A\\ -\frac{1}{n}A^{\top}&\frac{\rho\eta_{1}}{n^{2}}B\end{pmatrix},P_{4}=\begin{pmatrix}\rho^{-1}I&-\frac{1}{n}A\\ -\frac{1}{n}A^{\top}&\frac{\rho\eta_{2}}{n^{2}}I\end{pmatrix},$ where $A=X,B=\mathrm{diag}\left(X_{[1]}^{\top}X_{[1]},\ldots,X_{[K]}^{\top}X_{[K]}\right)$.
The generic update formula reveal structural similarity between Consensus-PD and Proximal-1-PD, and between LinConsensus-PD and Proximal-2-PD. When parameters $\eta_{1},\eta_{2},\tau$ are selected to ensure fastest convergence rate, i.e., $\eta_{1}=K$, $\tau=\max_{k\in[K]}\lambda_{\max}(X_{[k]}^{\top}X_{[k]})$, and $\eta_{2}=K\tau$, adjusting the step size of Consensus-type algorithms and Proximal-type algorithms such that $\beta K=\rho^{-1}$ will lead to the following compact forms of $P_{3}$ and $P_{4}$:
$P_{3}=\begin{pmatrix}\beta KI&-\frac{1}{n}A\\ -\frac{1}{n}A&\frac{\beta^{-1}}{n^{2}}B\end{pmatrix}\qquad P_{4}=\begin{pmatrix}\beta KI&-\frac{1}{n}A\\ -\frac{1}{n}A&\frac{\tau\beta^{-1}}{n^{2}}I\end{pmatrix}$.
It indicates $P_{3}$ and $P_{1}$ are the same, but the off-diagonal blocks flip the sign. $P_{2}$ and $P_{4}$ are the same, but the off-diagonal blocks flip the sign. However, how such connection relates to the convergence properties of each pair of algorithms remains unclear, and we will leave it for future study.
4 A Unified Proof of Convergence
@@
Representing the four ADMM-based primal-dual update forms (4)–(7) into the generic update rules also provides a unified and straightforward approach of their convergence analysis. Using the convergence analysis framework of He & Yuan (2012); Lu & Yang (2023), we establish an $O(1/T)$ ergodic rates of these algorithms. Theorem 4.1 characterizes the $O(1/T)$ ergodic convergence rate under exact updates.
Theorem 4.1.
@@
Let $\left\{z^{(t)}=(w^{(t)},v^{(t)})\right\}_{t=0}^{\infty}$ be the sequence generated by the generic update rule (9) with a positive semi-definite matrix $P$ and the initial point $z^{(0)}=(w^{(0)},v^{(0)})$. For any $z=(w,v)$, the following inequality holds:
@@
By selecting $z=(w^{*},v^{*})$, the optimal solutions to (1) and (2), Theorem 4.1 indates that $L(\bar{w}^{(T)};v^{*})-L(w^{*};\bar{v}^{(T)})$ converges to zero, which implies that $\bar{z}^{(T)}$ converges to the optimal solution with rate $O(1/T)$.
Theorem 4.1 guarantees that all starting points $z^{(0)}\in B_{P}(z^{*},M)=\{z:\left\lVert z-z^{*}\right\rVert_{P}<M\}$ converges with rate $M/2T$. The volume of $B_{P}(z^{*},1)$ is thereby an indicator for the overall convergence rate associated with a generic update rule with $P$, which proportional to $(\det P)^{-1}$. Clearly, $\det P_{1}=\det P_{3}$ and $\det P_{2}=\det P_{4}$, so the overall size of the bound provided by Theorem 4.1 are the same for Consensus-PD and Proximal-1-PD, and the same for LinConsensus-PD and Proximal-2-PD. Similar convergence behavior for both pairs of update rules are further validated in our experiment studies.
All $v$-steps of the updates (4)–(7) solve a minimization problem which would rely on an inner loop, in case no closed-form solution is available. We present a unified proof of convergence for (4)–(7) which addresses the inexact updates, based on the technique of Lu & Yang (2023).
Theorem 4.2.
Let $P$ be positive definite, and $z^{*}=(w^{*},v^{*})$ be the optimal solution of the saddle point problem (8). If the sequence $\{z^{(t)}\}$ satisfies
$P(z^{(t)}-z^{(t+1)})+\epsilon^{(t+1)}\in\mathcal{F}(z^{(t+1)})$ (10)
with $\sum_{t=1}^{\infty}\|\epsilon^{(t)}\|_{2}<\infty,$ then there exists a constant $D<\infty$ such that
$\sup_{t}\|z^{*}-z^{(t)}\|\leq D$
and
$L(\bar{w}^{(T)};v^{*})-L(w^{*};\bar{v}^{(T)})\leq\\ \frac{\|z^{*}-z^{(0)}\|_{P}^{2}}{2T}+\frac{D\sum_{t=1}^{T}\|\epsilon^{(t)}\|_{2}}{T}.$
@@
5 Experiment
In this section, we perform three experiments to test the performance of the primal-dual update rules under different parameter settings. The individual goals of the three experiments are as follows:
  • In experiment 1, we study how the the tuning parameters affect each algorithm, to verify our suggestions on tuning parameter selection.
  • In, experiment 2, we evaluate the performance of the five update rules using synthetic data for Lasso and Ridge regression tasks.
  • In experiment 3, we continue to evaluate the performance of the five update rules on three real-world binary classification tasks employing SVM.
Performance Metric. The relative gap difference is used to evaluate the performance of the algorithms in all experiments. It is defined as
$\epsilon_{G}:=\frac{|\mathcal{P}(w)+\mathcal{D}(v)|}{1+|\mathcal{P}(w)|+|\mathcal{D}(v)|},$
where $\mathcal{P}(w)$ and $\mathcal{D}(v)$ are calculated from (1) and (2) based on the $v$-step and $w$-step updates of each iteration $t$. Note that under our definition of the dual problem (2) the numerator gives the dual gap. The algorithm for which $\epsilon_{G}\to 0$ at a higher rate performs better.
5.1 Experiment 1
Experiment 1 aims to verify the effect of $\eta_{1},\eta_{2}$, $\tau$ for the Proximal-1-PD, Proximal-2-PD, and LinConsensus-PD update rules, and compare them with the effect of the step sizes $\rho$ and $\beta$.
Datasets. We used a1a dataset in the LibSVM library(Chang & Lin, 2011): a1a, w8a, and real-sim. Details of this dataset, including the number of samples, features, and clients, is included in Table 1. In the problem, we evenly distribute the data to $K=10$ machines.
Method. We train the model using $\ell_{2}$-regularized SVM model with regularization parameter of $\lambda=1/n$. We tested the performance of three algorithms, where Proximal-1-PD is subject to the tuning parameters $\eta_{1}$ and $\rho$, Proximal-2-PD is subject to the tuning parameters $\eta_{2}$ and $\rho$, and LinConsensus-PD is subject to $\tau$ and $\beta$. In the experiments, we test each method through fixing one tuning parameter and setting multiple values for the other tuning parameter. We record the trajectory of the relative gap difference in 500 communication rounds.
Results. The trajectory of the gap are shown in Figure 2. The first column demonstrated the validity of the selection of $\eta_{1},\eta_{2}$ and $\tau$ for the three methods. The recommended values, denoted by light blue lines, ensure the convergence of the algorithm. When these values become smaller, the convergence speed increases only slightly (e.g., the green and purple line). When these values become too small, however, the algorithms may fail to converge. Second, we can see from the three figures in second column that the ADMM step size parameter $\rho$ and $\beta$ has significantly impacts the algorithms’ performance.

### v2

@@ 3.2 Distributed Proximal ADMM
The work [4] introduces an additional proximal term into the standard ADMM algorithm for solving $\min_{x,y}f(x)+g(y)$ subject to $Ax+By=b$ , which is referred to as generalized ADMM or proximal ADMM. Applying Algorithm 2 of the work [4], the proximal ADMM solving the dual problem (D) updates as follows:
$\begin{split}v^{(t+1)}&=\underset{v\in\mathbb{R}^{n}}{\arg\min}\;\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{P}_{k}}\ell^{*}_{i}(v_{i})-\langle w^{(t)},\frac{1}{n}Xv+u^{(t)}\rangle+\frac{\rho}{2}\Big\|\frac{1}{n}Xv+u^{(t)}\Big\|^{2}+\frac{1}{2}\|v-v^{(t)}\|_{Q}^{2},\\ u^{(t+1)}&=\underset{u\in\mathbb{R}^{d}}{\arg\min}\;g^{*}(u)-\langle w^{(t)},\frac{1}{n}Xv^{(t+1)}+u\rangle+\frac{\rho}{2}\Big\|\frac{1}{n}Xv^{(t+1)}+u\Big\|^{2},\\ w^{(t+1)}&=w^{(t)}-\rho\left(\frac{1}{n}Xv^{(t+1)}+u^{(t+1)}\right),\end{split}$
where $\rho>0$ is the tuning parameter and $Q$ is a positive semi-definite matrix. To enable parallel updates of $v$ across $K$ machines, the following proposition gives two positive semi-definite matrices $Q$ choices.
Proposition 2.
After eliminating the auxiliary variable $u$, the proximal ADMM updates can be simplified by choosing appropriate proximal matrices. In particular:
1. when $Q_{1}=\frac{\rho}{n^{2}}(\eta_{1}\,\text{diag}(X_{[1]}^{\top}X_{[1]},\ldots,X_{[K]}^{\top}X_{[K]})-X^{\top}X)$, the proximal ADMM simplifies to
$w^{(t)}=$ $\mathrm{prox}_{\rho g}\left(w^{(t-1)}-\frac{\rho}{n}Xv^{(t)}\right),$ (3)
$v^{(t+1)}_{[k]}$= $\underset{v_{[k]}\in\mathbb{R}^{n_{k}}}{\arg\min}\frac{1}{n}\sum_{i\in\mathcal{P}_{k}}\ell^{*}_{i}(v_{i})+\frac{\rho\eta_{1}}{2n^{2}}\Big\|v_{[k]}-v_{[k]}^{(t)}\Big\|_{X_{[k]}^{\top}X_{[k]}}^{2}-\frac{1}{n}\Big\langle X_{[k]}^{\top}\left(2w^{(t)}-w^{(t-1)}\right),v_{[k]}\Big\rangle,k\in[K],$
2. when $Q_{2}=\frac{\rho}{n^{2}}\left(\eta_{2}I-X^{\top}X\right)$, the proximal ADMM simplifies to:
$\begin{split}w^{(t)}&=\mathrm{prox}_{\rho g}\left(w^{(t-1)}-\frac{\rho}{n}Xv^{(t)}\right),\\ v_{[k]}^{(t+1)}&=\mathrm{prox}_{(n/\rho\eta_{2})\ell^{*}_{[k]}}\left(v_{[k]}^{(t)}+\frac{n}{\rho\eta_{2}}X_{[k]}^{\top}\left(2w^{(t)}-w^{(t-1)}\right)\right),k\in[K].\end{split}$ (4)
The update rule of (3) and (4) are named Proximal-1-PD and Proximal-2-PD, respectively. The distributed proximal ADMM algorithms of either $Q$ are guaranteed to converge for any $\rho>0$. The following lemma provides a reliable choice for selecting the tuning parameters $\eta_{1}$ and $\eta_{2}$ that ensures $Q_{1}$ and $Q_{2}$ to be positive semidefinite, satisfying the requirement of [4].
Lemma 1.
@@
The minimal $\eta_{2}$ to let $Q_{2}\succeq 0$ is $\lambda_{\max}\left(X^{\top}X\right)$. However, this choice is practically infeasible in a distributed learning setup, as it requires the aggregation of samples from all machines.
3.3 CoCoA with Ridge Penalty
Unlike the aforementioned methods, which are applicable to general regularized ERM problems, the CoCoA framework was originally proposed to solve the dual of the $\ell_{2}$-regularized problem. Specifically, when the regularization term is the ridge penalty $g(w)=\frac{\lambda}{2}\|w\|_{2}^{2}$, CoCoA performs the following updates:
$\begin{split}\tilde{v}^{(t)}&=\arg\min_{v\in\mathbb{R}^{n}}\;\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{P}_{k}}\ell_{i}^{*}(v_{i})+\frac{1}{n^{2}\lambda}\left\langle X^{\top}Xv^{(t)},v\right\rangle+\frac{\sigma}{2n^{2}\lambda}\sum_{k=1}^{K}\left\|v_{[k]}-v_{[k]}^{(t)}\right\|^{2}_{X_{[k]}^{\top}X_{[k]}},\\ v^{(t+1)}&=v^{(t)}+\gamma\left(\tilde{v}^{(t)}-v^{(t)}\right),\quad w^{(t+1)}=-\frac{1}{n\lambda}\sum_{k=1}^{K}X_{[k]}v_{[k]}^{(t+1)},\end{split}$
where the $w$-step recovers the primal variable from the dual via the KKT conditions, and the $v$-step aims to reduce the dual objective $\mathcal{D}(v)$. The parameters $\sigma$ and $\gamma$ control the approximation quality of the dual subproblem and the update aggressiveness, respectively. It has been shown in [20] that setting $\gamma=1$ and $\sigma=K$ yields the fastest guaranteed convergence.
Under these parameter choices, CoCoA simplifies to the following updates involving iterative primal and dual variables $w$ and $v$, which we refer to as CoCoA-PD:
$w^{(t)}$ $=-\frac{1}{n\lambda}\sum_{k=1}^{K}X_{[k]}v^{(t)}_{[k]}$, (5)
$v^{(t+1)}_{[k]}$ $=\arg\min_{v_{[k]}\in\mathbb{R}^{n_{k}}}\;\frac{1}{n}\sum_{i\in\mathcal{P}_{k}}\ell_{i}^{*}(v_{i})-\frac{1}{n}\left\langle X_{[k]}^{\top}w^{(t)},v_{[k]}\right\rangle+\frac{K}{2n^{2}\lambda}\left\|v_{[k]}-v_{[k]}^{(t)}\right\|^{2}_{X_{[k]}^{\top}X_{[k]}},\quad k\in[K].$
@@ 3.4 Summary
The five algorithms—Consensus-PD, LinConsensus-PD, Proximal-1-PD, Proximal-2-PD, and CoCoA-PD (Equations (1)–(5))—can all be cast into a unified primal-dual update framework. This unified view allows us to analyze the structural connections among the algorithms and to develop a common convergence analysis, as presented in Section 5.
In all algorithms, the update of the dual block $v_{[k]}$ involves applying the proximal operator $\mathrm{prox}_{\ell_{[k]}^{*}}(\cdot)$ or solving a regularized quadratic problem on a linear combination of the current primal variable $w^{(t)}$ and the previous dual variable $v_{[k]}^{(t)}$. Similarly, the update of the primal variable $w$ involves a linear combination of the current iterate $w^{(t)}$, the current messages $X_{[k]}v_{[k]}^{(t+1)}$ received from individual machines, and the previous messages $X_{[k]}v_{[k]}^{(t)}$.
An immediate advantage of using the unified primal-dual update formulation, rather than the original algorithm-specific forms, is that it enables efficient evaluation of the duality gap, which provides a bound on the objective error. The duality gap can be computed by substituting the current iterates $\{v_{[k]}^{(t)}\}_{k=1}^{K}$ and $w^{(t)}$ into the primal objective (P) and the dual objective (D), respectively.
@@
We summarize the selection of tuning parameters in the five algorithms mentioned above. With fixed optimal parameters $\sigma=K$ and $\gamma=1$ in the CoCoA algorithm [20], CoCoA-PD does not have tuning parameters. The optimal selection of parameters $\eta_{1},\eta_{2}$ in Proximal-1-PD, Proximal-2-PD, and $\tau$ in LinConsensus-PD are given in this article and confirmed by the experiments. The step sizes $\beta$ of Consensus-PD and LinConsensus-PD, and the step size $\rho$ of the Proximal-1-PD and Proximal-2-PD significantly affect the convergence speed [1] and should be tuned in a case-specific manner, as validated in our experiments (See Section 6).
4 Connections Among Existing Algorithms
@@
Corollary 1 (Equivalence of CoCoA-PD and Proximal-1-PD).
For $\ell_{2}$-regularized ERM problems with $g(w)=\lambda\|w\|_{2}^{2}$, the update rules in (3) and (5) produce identical dual iterates $v^{(t)}$ when $\rho=\lambda^{-1}$ and the algorithms are initialized identically.
@@
It is also interesting to see if the connection between CoCoA-PD and Prixmal-1-PD can be extended to other CoCoA variants with general penalty $g$ [20]. To this question, we give a negative answer, because the connection in Corollary 1 relies on the same quadratic structure of the ridge penalty in CoCoA and the augmented Lagrangian in the Proximal ADMM algorithm.
@@
Consensus ADMM and Proximal ADMM.
For Consensus-PD and Proximal-1-PD, observe that the saddle-point formulation satisfies
$\min_{w\in\mathbb{R}^{d}}\max_{v\in\mathbb{R}^{n}}L(w;v)=\max_{w\in\mathbb{R}^{d}}\min_{v\in\mathbb{R}^{n}}(-L(w;v)).$
Hence, Proximal-1-PD can be interpreted as applying Consensus-PD to the equivalent saddle-point problem with negated objective L(w;v)-L(w;v). This equivalence is formalized in the following corollary.
Corollary 2 (Equivalence of Consensus ADMM and Proximal ADMM).
Assume identical initialization and augmented Lagrangian parameters satisfying $\beta K=\rho^{-1}$ in Consensus ADMM and Proximal ADMM. Then, we have (1) Consensus-PD is equivalent to Proximal-1-PD when $\eta_{1}=K$, (2) LinConsensus-PD is equivalent to Proximal-2-PD when $\eta_{2}=K\tau$.
Combining Corollaries 1 and 2, we observe that CoCoA variants arise as special cases of consensus ADMM under specific parameter settings, challenging the conclusion in [20] that CoCoA is fundamentally distinct.
5 Theoretical Analysis
@@
Representing the four ADMM-based primal-dual update forms (1)–(4) into the generic update rules also provides a unified and straightforward approach of their convergence analysis. Using the convergence analysis framework of [8, 15], we establish an $O(1/T)$ ergodic rates of these algorithms. To proceed, we first show that each algorithm can be viewed as [15] applied to the Lagrangian saddle-point problem, as formalized below.
Lemma 2.
Let $z=(w,v)$ denote the concatenated primal and dual variables, and define the monotone operator
$\mathcal{F}(z)=\begin{pmatrix}\partial_{w}L(w,v)\\ -\partial_{v}L(w,v)\end{pmatrix}$,
where $L(w,v)$ is the Lagrangian of the saddle-point problem. Then, the update rules of the algorithms (1), (2), (3), and (4)can all be written in the generic proximal form:
$P(z^{(t)}-z^{(t+1)})\in\mathcal{F}(z^{(t+1)}),$
with the corresponding matrix $P$ specified as follows:
$P_{1}=\begin{pmatrix}\beta KI&\frac{1}{n}A\\ \frac{1}{n}A^{\top}&\frac{1}{n^{2}\beta}B\end{pmatrix},P_{2}=\begin{pmatrix}\beta KI&\frac{1}{n}A\\ \frac{1}{n}A^{\top}&\frac{\tau}{n^{2}\beta}I\end{pmatrix},P_{3}=\begin{pmatrix}\rho^{-1}I&-\frac{1}{n}A\\ -\frac{1}{n}A^{\top}&\frac{\rho\eta_{1}}{n^{2}}B\end{pmatrix},P_{4}=\begin{pmatrix}\rho^{-1}I&-\frac{1}{n}A\\ -\frac{1}{n}A^{\top}&\frac{\rho\eta_{2}}{n^{2}}I\end{pmatrix}$
where $A=X$, and $B=\mathrm{diag}\left(X_{[1]}^{\top}X_{[1]},\ldots,X_{[K]}^{\top}X_{[K]}\right)$.
As the algorithm-specific matrices $P_{1},\ldots,P_{4}$ are positive semidefinite, the convergence analysis can be conducted within the standard framework of generalized PPMs [15, 14]. Specifically, Theorem 1 characterizes the convergence behavior of the general distributed primal–dual algorithmic framework.
Theorem 1.
@@
Let $\left\{z^{(t)}=(w^{(t)},v^{(t)})\right\}_{t=0}^{\infty}$ be the sequence generated by the generic update rule in Lemma 2 with a positive semi-definite matrix $P$ and the initial point $z^{(0)}=(w^{(0)},v^{(0)})$. For any $z=(w,v)$, the following inequality holds:
@@
By selecting $z=(w^{*},v^{*})$, the optimal solutions to (P) and (D), Theorem 1 indicates that $L(\bar{w}^{(T)};v^{*})-L(w^{*};\bar{v}^{(T)})$ converges to zero, which implies that $\bar{z}^{(T)}$ converges to the optimal solution with rate $O(1/T)$. In practice, all $v$-steps of the updates (1)–(4) solve a minimization problem which would rely on an inner loop, in case no closed-form solution is available. We present a unified proof of convergence for (1)–(4) which addresses the inexact updates, based on the technique of [15].
Theorem 2.
Let $P$ be positive definite, and $z^{*}=(w^{*},v^{*})$ be the optimal solution of the saddle point problem (SP). If the sequence $\{z^{(t)}\}$ satisfies $P(z^{(t)}-z^{(t+1)})+\epsilon^{(t+1)}\in\mathcal{F}(z^{(t+1)})$ with $\sum_{t=1}^{\infty}\|\epsilon^{(t)}\|_{2}<\infty,$ then there exists a constant $D<\infty$ such that $\sup_{t}\|z^{*}-z^{(t)}\|\leq D$and
$L(\bar{w}^{(T)};v^{*})-L(w^{*};\bar{v}^{(T)})\leq\frac{\|z^{*}-z^{(0)}\|_{P}^{2}}{2T}+\frac{D\sum_{t=1}^{T}\|\epsilon^{(t)}\|_{2}}{T}.$
@@
6 Experiments
In this section, we perform experiments to test the performance of the primal-dual update rules under different parameter settings. Specifically, we first studying how the the tuning parameters affect each algorithm, to verify our suggestions on tuning parameter selection. Then, we evaluate the performance of the five update rules using synthetic data for Lasso and Ridge regression tasks, which also verify the equivalency results in Section 4. Additional evaluations on the performance of the five update rules on three real-world binary classification tasks employing SVM are included in Appendix D.2. All experiments are conducted on the Dell Latitude 7450 Laptop, and each can be finished within 6 hours. The codes are available in supplementary material.
Experiment 1. We aim to verify the effect of $\eta_{1},\eta_{2}$, $\tau$ for the Proximal-1-PD, Proximal-2-PD, and LinConsensus-PD update rules, and compare them with the effect of the step sizes $\rho$ and $\beta$. We used a1a dataset in the LibSVM library[2]: a1a, w8a, and real-sim. Details of this dataset, including the number of samples, features, and clients, are included in Table 1. In the problem, we evenly distribute the data into $K=10$ machines. We train the model using $\ell_{2}$-regularized SVM model with regularization parameter of $\lambda=1/n$. We tested the performance of three algorithms, where Proximal-1-PD is subject to the tuning parameters $\eta_{1}$ and $\rho$, Proximal-2-PD is subject to the tuning parameters $\eta_{2}$ and $\rho$, and LinConsensus-PD is subject to $\tau$ and $\beta$. In the experiments, we test each method through fixing one tuning parameter and setting multiple values for the other tuning parameter. We record the trajectory of the relative gap difference in 500 communication rounds.
Table 1: Description of the datasets.
Dataset $n$ $d$ $K$
a1a 1605 119 10
w8a 49749 300 60
real-sim 72309 20958 100
Results. The trajectory of the gap are shown in Figure 2. The first row demonstrated the validity of the selection of $\eta_{1},\eta_{2}$ and $\tau$ for the three methods. The recommended values, denoted by light blue lines, ensure the convergence of the algorithm. When these values become smaller, the convergence speed increases only slightly (e.g., the green and purple line). When these values become too small, however, the algorithms may fail to converge. Second, we can see from the three figures in second row that the ADMM step size parameter $\rho$ and $\beta$ has significantly impacts the algorithms’ performance.